OpenStudy (anonymous):

find y' if 3xy^2 + y^3=e^(x+y)

OpenStudy (anonymous):

y' = [ e^(x+y) - 3y^2]/[6xy + 3y^2 - e^(x+y)]

OpenStudy (anonymous):

Be careful....d/dx of (3xy^2 + y^3) = 3y^2 +(3y^2)dy/dx d/dx(e^(x+y) = e^(x+y) times d/dx(x+y) = e^(x+y) OK so far? Can you work out the rest?